OpenStudy (anonymous):

Can someone help me please A general exponential demand function has the form q = Ae^−bp (A and b nonzero constants). (a) Obtain a formula for the price elasticity E of demand at a unit price of p. (b) Obtain a formula for the price p that maximizes revenue.

OpenStudy (anonymous):

elasticity=%(change in quantity)/%(change in quantity)

OpenStudy (anonymous):

when you start to calculate change in quantity=dq change in price =dp

OpenStudy (anonymous):

so %change in quantity=100*(dq/q) %change in price=100*(dp/p)

OpenStudy (anonymous):

therefore =>e=(100*dq/q)/100*(dp/p) e=(dq/q)*(p/dp) e=(dq/dp)*(p/q)

OpenStudy (anonymous):

now q=Ae^(-bp) dq/dp=A*-b*e^(-bp) e=p/q*A*-b*e^(-bp) e=p/Ae^(-bp)*{A*-b*e^(-bp)} e= -pb

OpenStudy (anonymous):

Revenue(r) =pq r=p*Ae^(−bp) dr/dp=Ae^(−bp)-pAbe^(−bp) dr/dp=Ae^(−bp)*{1-pb) for max revenue dr/dp=0 gives1-pb=0 gives p=1/b unite price that maximises revenue=1/b note that at this point |e|=1